A manufacturer of plasma television estimates that the average family uses its television 2,000 hours per year. The exponentially distributed mean time to failure for these televisions has been estimated to be 15,00 hours. The company firmly believes it is important to provide a reasonable warranty period for the television it sells; however, it realizes it would not be economically viable to service any more than 2% of the units it sells. What would be an appropriate warranty period ( expressed in terms a customers would understand)? Show all work organized!
Answer #1

here as we know that p th percentile =-ln(1-p)*\beta

therefore corresponding warranty period =-ln(1-0.02)*1500 =30.3040 Hours